OpenStudy (anonymous):

Linear algebra: Determine the basis of span{ [1,-1,1], [1,0,2], [1,-2,0], [1,1,-1]

OpenStudy (anonymous):

After row reduction i got: 1 0 0 -4 0 1 0 7/2 0 0 1 3/2 and now?

OpenStudy (anonymous):

um im pretty sure that you can just take th first three vectors and say they are a basis becuase hte 4th vector is just a linear combination of the 3

OpenStudy (anonymous):

X1 = 4 X4 X2 = 7/2 X4 X3 = 3/2 X4 X4 = free X4(4,7/2,3/2,1) ?

OpenStudy (anonymous):

{[1,-1,1], [1,0,2], [1,-2,0]} spans the same space as {[1,-1,1], [1,0,2], [1,-2,0], [1,1,-1]} and teh former is linearly independent so its a basis

OpenStudy (jamesj):

Or in other words, the four vectors span all of R^3 and hence {[1,0,0], [0 1 0], [0 0 1]} is also a basis.

OpenStudy (anonymous):

so x1, x2 and x3 span the basis

OpenStudy (jamesj):

x1, x2, x3 is a basis for the span

OpenStudy (anonymous):

x1 |4 | x = x2 = x4 |7/2| x3 |3/2| x4 |1 |

OpenStudy (jamesj):

No, that can't be right because if it were, the sub-space that is the span of those vectors would be one dimensional. How many dimensions does the span of { [1,-1,1], [1,0,2], [1,-2,0], [1,1,-1] } have?

OpenStudy (anonymous):

4

OpenStudy (jamesj):

Haven't you just shown that the first three vectors are linearly independent? So at least 3. Now 4 can't be right either because R^3 only has 3 dimensions. And it isn't the case that all four vectors are linearly independent.

OpenStudy (anonymous):

owh shoot... i ment 3 because i have 3 pivots

OpenStudy (jamesj):

Hence the span is three dimensional and any basis will have three basis vectors.

OpenStudy (jamesj):

the vectors [1,-1,1], [1,0,2], [1,-2,0] will do. But as this span is also R^3, the basis [1 0 0], [0 1 0], [0 0 01] also works.

OpenStudy (anonymous):

ahh oke.. thanks for the explanation :)